Harrison Water Sports has two retail outlets: Seattle and Portland. The Seattle store does 60 percent of the total sales in a year. Further analysis indicates that if a sale is made in Seattle, there is 0.40 probability this is a sale of boat accessories, while if a sale is made at the Portland store this probability is 0.20. If Harrison Water Sports knows that a boat accessory has been sold, what is the probability this sale has occurred in Seattle:

a.

0.24

b.

2/3

c.

3/4

d.

None of the above
Answer #1

We use Bayes Theorem

Let S and P denote the events of sale in Seattle and Portland.

Let B denote a sale of boat accessory.

Given,

P(S) = 60% = 0.6

P(P) = 1- 0.6 = 0.4

P(B/S) = 0.4

P(B/P) = 0.2

P(S)P(BS) P(SB)- PlP) * P( BIP) + P(S) * P(BIS)

0.6 0.4 3/40.75 0.4 * 0.20.6 0.4

Ans is c) 3/4